Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 9/5/2017

Location : CR21 at 14:00 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

Agenda

  • Actions
    • Macwaters - Re-make dodgy compressor2 remote cable connector - stands
    • Macwaters - Remove old overheated CSW71 to R9. Speak to Sumitomo on April 5th regards how to repair. - in progress
    • Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress
    • Macwaters - Report on status of FC psu repair (stands)
    • Hall Manager - water system change-over procedure (stands)
    • EOverton - Status of miceraid5 disk cleanup (stands)
    • JBoehm & Macwaters - get the West Wall compressors correctly labelled, as various have been re-purposed (stands)
    • The Picologger has been disconnected. SBoyd to check with CWhyte whether we want it or not
      • Contacted Colin/Josef : Josef will meet with John Cobb on Friday to discuss this in the context of ramping the FC

AOB:

Minutes

  • Actions
    • Macwaters - Re-make dodgy compressor2 remote cable connector - stands
    • Macwaters - Prepare new F70 compressors for use. Remove old overheated CSW71 to R9. Speak to Sumitomo on April 5th regards how to repair. - in progress
    • Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress
    • Macwaters - Report on status of FC psu repair (stands)
      • TH chasing Thurlby-Thandar. Cause of failure is that there may still have been a few mA in the circuit when it was powered off with contactors still closed
      • The spare is installed and has been run into a short
    • Hall Manager - water system change-over procedure (stands)
    • EOverton - Status of miceraid5 disk cleanup (stands)
    • MCourthold - show JBoehm how to use the Picologger for quench data. Craig can help. (on hold)
      • The Picologger has been disconnected. SBoyd to check with CWhyte whether we want it or not
      • JC has shown JB how to use the Picologger which has been re-connected to the FC quench system
    • JBoehm & Macwaters - get the West Wall compressors correctly labelled, as various have been re-purposed (stands) Use the number before the # symbol - this is expected to agree with the Alarm Handler
    • CWhyte to contact DL to clarify their Hall Task List items
  • New Actions
    • CWhyte to request survey of TOF1
  • DC Update:
    Activities mainly LH2 work and Tracker cryo saga: seems to be a lot of air getting in. Lots of RGA scans
    • Tracker Issue
      Cryostat 2: followed the official procedure for warm-up and cool-down. Have lost 12% of channels on one cassette and 27% on the other. Thought to be due to aging seals.
      Pump/purged Cryo 3 yesterday, today see 13% N2. Think there's an air leak in. Try a pump/purge excluding the cassette space to see if the leak is at the cassette or the plumbing.
      Also had a "Low Lid Temperature" alarm and a backing pump fail on Cryo 2 today.
    • LH2 Update
      The LH2 Quench test with Neon will be in June (no CC magnets this cycle).
      Hope to be ready to start cooling today - final vacuum and leak-checking - and start cool-down tomorrow. Still need to fix the quench line before the real LH2. FC will stay in position this cycle.
  • MOM Update
    Expert start-up yesterday: The wrong absorber got written into the CDB, else OK (note have no Trackers).
    AKurup wants to try new ALH tonight.
    Beambump tune today: search Hall at 4 - with LH2 and Tracker work, better to postpone till tomorrow. We need ISIS BLM info to be getting into EPICS.
    Daily Run Meetings will be (in ISIS coffee lounge) at 2 p.m. so that we know how the Hall work is going that day.
    Next tranche of Shifter training will be on Monday.
  • AOB
    TOF1 has been replaced in position and the plastic sheet removed: it will need a survey.

Updated by Nebrensky, Henry about 6 years ago · 3 revisions